Sharp Inc., a wholly Ghanaian owned company specializes in the production of hand 

sanitizers branded as Quin. The company has branches in all the regions in Ghana with its 

head office located at Cape Coast. It employs 1,200 people and pays an average of GHȼ

550.00 as income tax to the government of Ghana on a monthly basis. Suppose that seven 

thousand, eight hundred and twenty (7,820) units of Quin are produced and supplied by Sharp 

Inc. but the quantity demanded for Quin is eight thousand (8,000) units. Ceteris Paribus, a 

GHȼ 20.00 change in the price of Quin results in a change in quantity demanded for and 

supplied of 4 and 5 units, respectively. [Hint: change in X due to change in Y is given as ]. 

a) Determine the equilibrium price and equilibrium quantities of Quin.

b) Infer the elasticity of demand for Quin from your solution.

 



1
Expert's answer
2021-06-12T10:43:36-0400

q=7820


"E=\frac{\delta Q}{\delta p}=\frac{5-4}{20}=0.05"